# From MRI to SARI: Parametric survey of non-axisymmetric accretion disk instabilities

### v1

@@ 5.4.3 Global vs local approach
The global eigenvalue problem as in Keppens et al. (2002) has often been in the shadows of the local WKB approximation because of the involved analytical methods required. However, the interaction of nearby singularities of the governing ODE is exactly what is responsible for the special properties of non-axisymmetric modes. We briefly show in this section that local dispersion relations are unable to correctly predict both discrete and quasi-continuum SARI modes. In the Appendix B, we show how such a dispersion relation can be derived from the global formulation using a WKB approach (Keppens et al., 2002; Blokland et al., 2005), and how some recent dispersion relations obtained directly from the linearised MHD equations also follow from the global formulation (Das et al., 2018; Ebrahimi & Pharr, 2022).
@@
Solving a local WKB dispersion relation requires a radius $r$ where all equilibrium parameters should be evaluated, and a radial wavenumber $q$ in addition to $k$ and $m$. Blokland et al. (2005) showed that a sixth-order local dispersion relation can exactly replicate MRI eigenfrequencies if $r$ and $q$ are derived self-consistently from the peak radius and radial wavelength of an exact eigenfunction of the global formulation. We now evaluate this dispersion relation over the whole radial domain to obtain the growth rates of the most unstable modes, and consider several values of $q$. In Fig. 10, a range of predicted eigenfrequencies from the sixth-order dispersion relation (B4) is plotted for three fixed values of $q$ (orange). For the most unstable inner SARI, this leads to a good approximation ($q\approx 50$ is the value obtained from the eigenfunction). The same cannot be said for the quasi-continuum SARI modes: since the eigenfunctions for modes with growth rate $\nu$ are very similar in shape but shifted horizontally according to the local Doppler resonance, a constant value of $q$ would have to correspond with a nearly horizontal band of quasi-continuum modes. However, as Fig. 10 shows, the maximal growth rates are quickly decreasing towards the outer edge of the disk and the dispersion relation predicts a stabilisation of all modes for all three radial wavenumbers. Note that the larger wavenumbers $q$, which should correspond to radially localised modes, perform worse in predicting ranges of unstable modes, and even start missing out the entire Doppler range. Indeed, also in the local dispersion relation of Terquem & Papaloizou (1996), the MRI is stabilised for $q\gg k$. The WKB analysis assumes radially oscillating solutions over the whole domain, whereas SARI eigenfunctions are radially localised. MRI eigenfunctions can fill the entire radial domain and hence are better suited for such an approximation. In particular, the dispersion relation gives no correct predictions for the outer SARI growth rates. The local WKB approximation is hence unable to replicate the complex interactions of the two singularities, which are crucial in the analysis for SARI modes. In contrast, the green line in Fig. 10 representing $\nu_{\text{max}}$ (9) obtained through global means, provides an upper bound for the discrete SARIs as well as the quasi-continuum modes, and correctly predicts the radial decrease of the maximum growth rates.
@@ 5.5 Extremely localised modes in global spectra
In Sect. 3 and 4, we have shown typical MRI and SARI spectra (Fig. 1), as well as unstable 2D continua filled with quasi-modes (Fig. 3). We then proceeded in Sect. 5 to calculate growth rates of the most unstable MRI and SARIs over a range of parameters, showing how the inner SARIs usually dominate over the outer SARIs and how modes can be stabilised by locally vanishing $\omega_{A}$. However, not just the fastest-growing modes are of interest in this study: the presence of quasi-continua provides a source of unstable modes with localised eigenfunctions all over the disk, as shown in Sect. 4. We set forth to calculate a quasi-continuum of large-$m$ SARIs that are extremely localised (but global in nature).
In principle, a SARI quasi-continuum can occur whenever $\Omega_{A}^{+}$ and $\Omega_{A}^{-}$ overlap, but the height that it might reach can be severely limited, requiring a high resolution ($n_{\text{grid}}\geq 1000$) to adequately resolve its edges. This depends on the ratio $|k|/|m|$, which appears in the exponent of exact quasi-continuum solutions (Goedbloed & Keppens, 2022). As this ratio gets large, the associated complementary energy greatly decreases and as a result, the quasi-continuum ‘grows’. These two conditions are summarised in Eq. (91) of Goedbloed & Keppens (2022), which we repeat here,
@@
We use Eq. (12) and the results of Sect. 5 to obtain a SARI quasi-continuum for a high azimuthal mode number $m=-10$. The two conditions need to be balanced: the ratio $|k/m|$ must be large, so a $k\gg|m|$ is needed – for Fig. 3, $k/|m|=50$ – but at the same time the instability has a cut-off with $k$ when the magnetic tension of the vertical field component stabilises the mode, as seen in Sect. 5.1. Additionally, the overlap between $\Omega_{A}^{\pm}$ strongly depends on $\epsilon^{-1}$, which needs to be large. A high-$m$ quasi-continuum hence prevails in weak fields, where large $k$ are allowed for instability. Alternatively, a stronger toroidal component moves $k_{\text{cut-off}}$ to larger values as shown in Sect. 5.2 and thus can also allow large $k$. An additional difficulty is that, for strongly overlapping $\Omega_{A}^{\pm}$, SARI resonances are closeby and the modes are radially very localised between those resonances, requiring a high resolution for resolving the eigenfunctions. All in all, we use the equilibrium parameters $\mu_{1}=10$, $\beta=10$, $\epsilon=0.01$, $\delta=1$ and $k=400$ so that $k/|m|=40$.
@@
Figure 11 shows the results of four QR-runs with increasing resolution that gradually find the true edges of a quasi-continuum for $m=-10$, overlaid on a quasi-continuum calculated with ROC at two values of $W_{\text{com}}$. Both the left and upper edges of the quasi-continuum are only fully resolved at a Legolas resolution of $n_{\text{grid}}=1000$. The latter has a value of $\nu\approx 0.2$, which is comparable to the results of Fig. 3. The scattered modes found within the quasi-continuum bounds are part of dense clouds of eigenvalues when applying Arnoldi-shift-invert to locations inside the 2D region (not shown). Furthermore, the eigenfunctions, shown in Panel (b), again resemble wave packages localised between the Alfvén resonances, which are now minimally separated (only $\sim 1\%$ of the radial domain!). Hence, the criteria from Sect. 4 to classify this Legolas spectrum as a quasi-continuum have been met. Using a modified version of ROC that enables shooting between the very closeby Alfvén resonances instead of over the whole domain, we see that the eigenfunctions oscillate heavily for $\nu\rightarrow 0$ and that they fill the whole region between the resonances. All together, we have modes that are highly localised in all directions: azimuthally with $m=-10$, longitudinally with $k=400$ and strongly oscillating radially in a small portion of the accretion disk, far away from the boundaries. Hence, these modes are truly localised, even though they are global in nature. It is hence not unthinkable that they have shearing box analogs, where only a tiny radial extent still produces overlapping continua that generate SARI-like modes (Matsumoto & Tajima, 1995; Noguchi et al., 2000).
@@
Figure 11: A large-$m$ ($-10$), large-$k$ ($400$) quasi-continuum with radially very localised eigenfunction $\Re(v_{r})$ (cyan) ($\omega=-7.62+0.09i$) as a result of the large overlap between $\Omega_{A}^{\pm}$. Increasing grid resolution finds edges that match with $W_{\text{com}}=10^{-12}$ (black). Equilibrium parameters: $\epsilon=0.01$, $\mu_{1}=10$, $\beta=10$, $\delta=1$.
@@ 6 Discussion and conclusions
The first part of this work considered the SARI quasi-continua discovered by Goedbloed & Keppens (2022). We independently confirmed their existence, noting that they are truly indistinguishable from true normal modes by any numerical code - and physical plasma, for that matter. Although the mathematical treatment of these modes is quite novel, hints of their existence were already present in earlier works: the non-axisymmetric spectra of Keppens et al. (2002) show large collections of modes that are reminiscent of the spectra in Fig. 2, where the QR-algorithm finds many scattered modes inside a quasi-continuum region. An earlier example are the modes excited in two-dimensional simulations by Terquem & Papaloizou (1996), which have some resemblance with the quasi-mode eigenfunctions and move at the local Doppler speed. More analysis is needed to determine whether they are truly like the wave packages seen in Sect. 4. In any case, it is clear that a global treatment of the problem is essential to capture these effects: the nearness of singularities in the overlapping continua lie at the origin of the problem, and these cannot be obtained through a local analysis. With the definition of quasi-modes having a tiny but finite complementary energy on a firm physical and mathematical basis, we presented how the quasi-modes can be identified using modern MHD spectroscopy with a tool like Legolas. The ROC and Legolas codes can be operated in tandem, so that the predictory value of the latter can be tested against the explicit $W_{\text{com}}$ calculations of the former, as was demonstrated in Fig. 11. It will be interesting to establish whether a shearing box equivalent of SARIs exists in our formalism, resembling the radially localised non-axisymmetric modes found by Matsumoto & Tajima (1995) and Noguchi et al. (2000). Global high-$m$ modes with eigenfunctions that are adequately localised could perhaps even survive in the small radial variation of a local box.
@@
In the second part, we performed a comprehensive study of instability growth rates in a wide range of equilibria, encompassing accretion disks with weak to strong, vertical to azimuthal, and subthermal to suprathermal magnetic fields. The stabilisation of modes at various values of $m$ was investigated for the most unstable inner and outer SARIs and the MRI. It was shown that stability is indeed governed by a condition (10) on the Alfvén frequency that follows from analytical theory, which was used to explain growth rates for weakly magnetised, thin disks. A recent upper bound for the growth rates (9) explained why the outer SARI growth rates behave like a scaled-down version of their inner equivalents. We also concluded that the $m<0$ counter-rotating SARIs play an important role in opening up more regimes for instability, significantly extending the range of unstable equilibria at high $k$, as shown in Fig. 4 and Fig. 7 (b). Our findings for non-axisymmetric modes and their implications are summarised below. In particular, the maximal values of the numerically calculated growth rates are $\sim 0.6\Omega_{1}$, in correspondence with the global analysis of Terquem & Papaloizou (1996), Begelman et al. (2022), while Balbus & Hawley (1992) find growth rates of only several percent of $\Omega_{1}$ in a local shearing sheet approximation. This again highlights the importance of using a global method when dealing with non-axisymmetric modes.
@@
Localisation – in a thin, weakly magnetised disk, the fastest-growing mode is found at intermediate to small wavelengths, with both $k_{\text{max}}$ and $k_{\text{cut-off}}$ increasing with increasingly negative $m$ as instability depends on $\omega_{A}$, a conclusion also drawn by Begelman et al. (2022). For equilibria with a sizeable toroidal field, the largest growth rates are found at large $k$ (Balbus & Hawley, 1992; Ogilvie & Pringle, 1996). This implies that the essence of the instability is still captured in this approximation of an infinitely long cylinder, which requires the assumption of small vertical wavelengths compared to the scale height, and the modes found represent modes that are localised around the midplane of a 3D accretion disk (Balbus & Hawley, 1998). In contrast, in models of accretion tori and 3D shearing boxes, unstable modes near the vertical edges of the disk are found (Balbus & Hawley, 1998), as well as strongest growth for the MRI at wavelengths about the size of the vertical scale height $H$ (Mamatsashvili et al., 2013).
@@
Additionally, we showed in Sect. 5.5 that large-$m$ modes have significant maximum growth rates, usually lower than the MRI but sometimes comparable. However, given the vast range of accessible wavelengths for non-axisymmetric instabilities, even those with growth close-to-optimal can dynamically outperform the axisymmetric MRI (Mamatsashvili et al., 2013). In any case, the obtained growth rates are an order of magnitude larger than those applicable to the shearing sheet model by Balbus & Hawley (1992), as was also remarked by Curry & Pudritz (1996). The possibility of a locally vanishing $\omega_{A}$ locally stabilises SARI modes. We also explicitly showed that quasi-continua do exist in configurations with large $m$ and $k$ (Fig. 11). Quasi-continuum SARI modes with such large wavenumbers are hence also very relevant to shearing box simulations, where the simulation domain is both radially, vertically and azimuthally localised (Lesaffre et al., 2009; Bai & Stone, 2013; Hirai et al., 2018). The quasi-modes resulting from these spectra can corotate anywhere in the radial domain, making them ideal candidates for instability in both local Cartesian and global simulations.
@@
In practice, however, large $k$ and $m$ modes can only be realised as long as the scales involved are above resistive and viscous length scales. The effect of these non-ideal terms wil generally be to quench the instability, in particular on smaller scales (Balbus & Hawley, 1991, 1998; Kitchatinov & Rüdiger, 2010). We plan to further investigate the effect of these diffusive terms on the MRI/SARI, and especially on the localised quasi-continuum SARI modes. Since the inclusion of non-ideal MHD effects breaks the up-down symmetry of MHD spectra (Goedbloed et al., 2019), it will be interesting to see how the quasi-continuum regions are modified in the structure of the spectra. We note that Legolas can analyse a number of non-ideal effects in combination, and thus can explore visco-resistive, or Hall and ambipolar effects on the ideal MHD spectrum.
@@
Magnetic field – We can conclude from Sect. 5.2 and 5.3 that a strong poloidal field inhibits large-$k$ instabilities, while a strong toroidal field inhibits large-$m>0$ instabilities: the magnetic tension produced by these small-wavelength modes would be too high (Balbus & Hawley, 1998). As a result, if the field is predominantly poloidal and strong, the MRI will be stabilised at small $k$ while negative $m$ modes can still have significant growth rates for large $k$ if $\omega_{A}$ satisfies criterion (10) (Begelman et al., 2022; Begelman & Armitage, 2023), as is also confirmed by Fig. 5. In particular, counter-rotating SARIs have one value of $m$ where the unstable range of field strengths is maximally extended, as on Fig. 7 (b). This implies that strong, poloidal fields could be unstable to extremely high-$m$ modes beyond the capabilities of our linear (and non-linear) codes: their eigenfunctions would be constrained to a tiny fraction of the domain. We also confirmed that for a purely toroidal field, a ‘complete breakdown of the most rapidly growing local modes’ (Balbus & Hawley, 1998) does indeed occur: $k_{\text{max}}\rightarrow\infty$ for purely toroidal field predicted by Ogilvie & Pringle (1996) quickly moves to much smaller values for even a seed poloidal field, as was noted in Fig. 6. Finally, in our exploration of the plasma-beta, we found that a dominant toroidal field has low growth rates for $\beta<1$. This was also found by Das et al. (2018) at intermediate field strengths, but they identified a new hybrid Alfvén/slow instability at stronger fields. It is unclear if a similar instability appears for our $m=10$ SARI in Case (a). According to van der Swaluw et al. (2005), these instabilities must be more MRI-like than convection-like, since the parameter range explored here corresponds to thin disks ($\sqrt{p_{1}}\leq 0.6$).
@@
Not only the fastest-growing modes are of importance for stability: the linear evolution is governed by an entire MHD spectrum of eigenmodes. We found that the outer SARI branch is usually stabilised first, implying that the quasi-continuum in between the discrete SARI branches can detach from the real axis and disappear before that, or continue to exist until the inner SARIs are also stabilised. It should be noted that we did not cover SARI quasi-continua for a dominant poloidal field, but these do exist. We confirmed this for $m=1$ at weak field strength. The Appendix covers the basic theory needed for the global formulation of the eigenvalue problem, from which a local dispersion relation can be derived. We showed the connection between the general dispersion relation from Blokland et al. (2005) and recent work by Das et al. (2018) and Ebrahimi & Pharr (2022). By comparing solutions to this dispersion relation with a SARI spectrum in Fig. 10, we highlighted that the local approximation is insufficient to describe outer and quasi-continuum SARI eigenmodes.
@@
NB is supported by an FWO Flanders fellowship titled “Magnetoseismology of accretion disks: towards the full spectrum for turbulent disks” - grant number 11J2622N. RK acknowledges support by the ERC Advanced Grant PROMINENT from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 833251 PROMINENT ERC-ADG 2018), the C1 project TRACESpace funded by KU Leuven and a FWO project G0B4521N.

### v2

@@ 5.4.3 Global vs local approach
The global eigenvalue problem as in Keppens et al. (2002) has often been in the shadow of the local WKB approximation because of the involved methods required for its analysis. However, the interaction of nearby singularities of the governing ODE is exactly what is responsible for the special properties of non-axisymmetric modes. We briefly show in this section that local dispersion relations are unable to correctly predict both discrete and quasi-continuum SARI modes. In Appendix B, we show how such a dispersion relation can be derived from the global formulation using a WKB approach (Keppens et al., 2002; Blokland et al., 2005), and how some recent dispersion relations obtained directly from the linearised MHD equations also follow from the global formulation (Das et al., 2018; Ebrahimi & Pharr, 2022).
@@
Solving a local WKB dispersion relation requires a radius $r$ where all equilibrium parameters should be evaluated, and a radial wavenumber $q$ in addition to $k$ and $m$. Blokland et al. (2005) showed that a sixth-order local dispersion relation can exactly replicate MRI eigenfrequencies if $r$ and $q$ are derived self-consistently from the peak radius and radial wavelength of an exact eigenfunction of the global formulation. We now evaluate this dispersion relation over the whole radial domain to obtain the growth rates of the most unstable modes, and consider several values of $q$. In Fig. 10, a range of predicted eigenfrequencies from the sixth-order dispersion relation (B4) is plotted for three fixed values of $q$ (orange). For the most unstable inner SARI, this leads to a good approximation ($q\approx 50$ is the value obtained from the eigenfunction). The same cannot be said for the quasi-continuum SARI modes: since the eigenfunctions for modes with growth rate $\nu$ are very similar in shape but shifted horizontally according to the local Doppler resonance, a constant value of $q$ would have to correspond with a nearly horizontal band of quasi-continuum modes. However, as Fig. 10 shows, the maximal growth rates are quickly decreasing towards the outer edge of the disk and the dispersion relation predicts a stabilisation of all modes for all three radial wavenumbers. Note that the larger wavenumbers $q$, which should correspond to radially localised modes, perform worse in predicting ranges of unstable modes, and even start missing out the entire Doppler range. Indeed, also in the local dispersion relation of Terquem & Papaloizou (1996), the MRI is stabilised for $q\gg k$. The WKB analysis assumes radially oscillating solutions over the whole domain, whereas SARI eigenfunctions are radially localised. MRI eigenfunctions can fill the entire radial domain and hence are better suited for such an approximation. In particular, the dispersion relation gives no correct predictions for the outer SARI growth rates. The local WKB approximation is hence unable to replicate the complex interactions of the two singularities, which are crucial in the analysis of SARI modes. In contrast, the green line in Fig. 10 representing $\nu_{\text{max}}$ (9) obtained through global means, provides an upper bound for the discrete SARIs as well as for the quasi-continuum modes, and correctly predicts the radial decrease of the maximum growth rates.
@@ 5.5 Extremely localised modes in global spectra
In Sect. 3 and 4, we have shown typical MRI and SARI spectra (Fig. 1), as well as unstable 2D continua filled with quasi-modes (Fig. 3). We then proceeded in Sect. 5 to calculate the growth rates of the most unstable MRI and SARIs over a range of parameters, showing how the inner SARIs usually dominate over the outer SARIs and how modes can be stabilised by locally vanishing $\omega_{A}$. However, not just the fastest-growing modes are of interest in this study: the presence of quasi-continua provides a source of unstable modes with localised eigenfunctions all over the disk, as shown in Sect. 4. We set forth to calculate a quasi-continuum of large-$m$ SARIs that are extremely localised (but global in nature).
In principle, a SARI quasi-continuum can occur whenever $\Omega_{A}^{+}$ and $\Omega_{A}^{-}$ overlap, but the height that it might reach can be severely limited, requiring a high resolution ($n_{\text{grid}}\geq 1000$) to adequately resolve its edges. This depends on the ratio $|k|/|m|$, which appears in the exponent of exact quasi-continuum solutions (Goedbloed & Keppens, 2022). As this ratio increases, the associated complementary energy greatly decreases and as a result, the quasi-continuum ‘grows’. These two conditions are summarised in Eq. (91) of Goedbloed & Keppens (2022), which we repeat here:
@@
We use Eq. (12) and the results of Sect. 5 to obtain a SARI quasi-continuum for a high azimuthal mode number $m=-10$. The two conditions need to be balanced: the ratio $|k/m|$ must be high, so a $k\gg|m|$ is needed – for Fig. 3, $k/|m|=50$ – but at the same time the instability has a cut-off with $k$ when the magnetic tension of the vertical field component stabilises the mode, as seen in Sect. 5.1. Additionally, the overlap between $\Omega_{A}^{\pm}$ strongly depends on $\epsilon^{-1}$, which needs to be large. A high-$m$ quasi-continuum hence prevails in weak fields, where large $k$ are allowed for instability. Alternatively, a stronger toroidal component moves $k_{\text{cut-off}}$ to larger values as shown in Sect. 5.2 and thus can also allow large $k$. An additional difficulty is that, for strongly overlapping $\Omega_{A}^{\pm}$, SARI resonances are close by and the modes are radially very localised between those resonances, requiring a high resolution for resolving the eigenfunctions. All in all, we use the equilibrium parameters $\mu_{1}=10$, $\beta=10$, $\epsilon=0.01$, $\delta=1$ and $k=400$ so that $k/|m|=40$.
@@
Figure 11 shows the results of four QR-runs with increasing resolution that gradually find the true edges of a quasi-continuum for $m=-10$, overlaid on a quasi-continuum calculated with ROC at two values of $W_{\text{com}}$. Both the left and upper edges of the quasi-continuum are only fully resolved at a Legolas resolution of $n_{\text{grid}}=1000$. The latter has a value of $\nu\approx 0.2$, which is comparable to the results of Fig. 3. The scattered modes found within the quasi-continuum bounds are part of dense clouds of eigenvalues when applying Arnoldi-shift-invert to locations inside the 2D region (not shown). Furthermore, the eigenfunctions, shown in Fig. 11 (b), again resemble wave packages localised between the Alfvén resonances, which are now minimally separated (by only $\sim 1\%$ of the radial domain!). Hence, the criteria from Sect. 4 to classify this Legolas spectrum as a quasi-continuum have been met. Using a modified version of ROC that enables shooting between the very close-by Alfvén resonances rather than over the whole domain, we see that the eigenfunctions oscillate heavily for $\nu\rightarrow 0$ and that they fill the whole region between the resonances. All together, we have modes that are highly localised in all directions: azimuthally with $m=-10$, longitudinally with $k=400$ and strongly oscillating radially in a small portion of the accretion disk, far away from the boundaries. Hence, these modes are truly localised, even though they are global in nature. It is hence not unthinkable that they have shearing box analogs, where only a tiny radial extent still produces overlapping continua that generate SARI-like modes (Matsumoto & Tajima, 1995; Noguchi et al., 2000).
@@
Figure 11: A large-$m$ ($-10$), large-$k$ ($400$) quasi-continuum with radially very localised eigenfunction $\Re(v_{r})$ (cyan) ($\omega=-7.62+0.09i$) as a result of the large overlap between $\Omega_{A}^{\pm}$. Increasing grid resolution finds edges that match $W_{\text{com}}=10^{-12}$ (black). Equilibrium parameters: $\epsilon=0.01$, $\mu_{1}=10$, $\beta=10$, $\delta=1$.
@@ 6 Discussion and conclusions
The first part of this work considered the SARI quasi-continua discovered by Goedbloed & Keppens (2022). We independently confirmed their existence, noting that they are truly indistinguishable from true normal modes by any numerical code - and physical plasma, for that matter. Although the mathematical treatment of these modes is quite novel, hints of their existence were already present in earlier works: the non-axisymmetric spectra of Keppens et al. (2002) show large collections of modes that are reminiscent of the spectra in Fig. 2, where the QR-algorithm finds many scattered modes inside a quasi-continuum region. An earlier example are the modes excited in two-dimensional simulations by Terquem & Papaloizou (1996), which have some resemblance with the quasi-mode eigenfunctions and move at the local Doppler speed. More analysis is needed to determine whether they are truly like the wave packages seen in Sect. 4. In any case, it is clear that a global treatment of the problem is essential to capture these effects: the nearness of singularities in the overlapping continua lies at the origin of the problem, and this cannot be obtained through a local analysis. With the definition of quasi-modes having a tiny but finite complementary energy on a firm physical and mathematical basis, we presented how quasi-modes can be identified using modern MHD spectroscopy with a tool like Legolas. The ROC and Legolas codes can be operated in tandem, so that the predictory value of the latter can be tested against the explicit $W_{\text{com}}$ calculations of the former, as was demonstrated in Fig. 11. It will be interesting to establish whether a shearing box equivalent of SARIs exists in our formalism, resembling the radially localised non-axisymmetric modes found by Matsumoto & Tajima (1995) and Noguchi et al. (2000). Global high-$m$ modes with eigenfunctions that are adequately localised could perhaps even survive in the small radial variation of a local box.
@@
In the second part, we performed a comprehensive study of instability growth rates in a wide range of equilibria, encompassing accretion disks with weak to strong, vertical to azimuthal, and subthermal to suprathermal magnetic fields. The stabilisation of modes at various values of $m$ was investigated for the most unstable inner and outer SARIs and the MRI. It was shown that stability is indeed governed by a condition (10) on the Alfvén frequency that follows from analytical theory, which was used to explain growth rates for weakly magnetised, thin disks. A recent upper bound for the growth rates (9) explained why the outer SARI growth rates behave like a scaled-down version of their inner equivalents. We also concluded that the $m<0$ counter-rotating SARIs play an important role in opening up more regimes for instability, significantly extending the range of unstable equilibria at high $k$, as shown in Fig. 4 and Fig. 7 (b). Our findings for non-axisymmetric modes and their implications are summarised below. In particular, the maximal values of the numerically calculated growth rates are $\sim 0.6\Omega_{1}$, in correspondence with the global analysis of Terquem & Papaloizou (1996) and Begelman et al. (2022), while Balbus & Hawley (1992) find growth rates of only several percent of $\Omega_{1}$ in a local shearing sheet approximation. This again highlights the importance of using a global method when dealing with non-axisymmetric modes.
@@
Localisation – in a thin, weakly magnetised disk, the fastest-growing mode is found at intermediate to small wavelengths, with both $k_{\text{max}}$ and $k_{\text{cut-off}}$ increasing with increasingly negative $m$ as instability depends on $\omega_{A}$, a conclusion also drawn by Begelman et al. (2022). For equilibria with a sizeable toroidal field, the highest growth rates are found at large $k$ (Balbus & Hawley, 1992; Ogilvie & Pringle, 1996). This implies that the essence of the instability is still captured in this approximation of an infinitely long cylinder, which requires the assumption of small vertical wavelengths compared to the scale height, and the modes found represent modes that are localised around the midplane of a 3D accretion disk (Balbus & Hawley, 1998). In contrast, in models of accretion tori and 3D shearing boxes, unstable modes near the vertical edges of the disk are found (Balbus & Hawley, 1998), as well as strongest growth for the MRI at wavelengths about the size of the vertical scale height $H$ (Mamatsashvili et al., 2013).
@@
Additionally, we showed in Sect. 5.5 that large-$m$ modes have significant maximum growth rates, usually lower than those of the MRI but sometimes comparable. However, given the vast range of accessible wavelengths for non-axisymmetric instabilities, even those with growth close-to-optimal can dynamically outperform the axisymmetric MRI (Mamatsashvili et al., 2013). In any case, the obtained growth rates are an order of magnitude higher than those applicable to the shearing sheet model by Balbus & Hawley (1992), as was also remarked by Curry & Pudritz (1996). The possibility of a locally vanishing $\omega_{A}$ locally stabilises SARI modes. We also explicitly showed that quasi-continua do exist in configurations with large $m$ and $k$ (Fig. 11). Quasi-continuum SARI modes with such large wavenumbers are hence also very relevant to shearing box simulations, where the simulation domain is radially, vertically, and azimuthally localised (Lesaffre et al., 2009; Bai & Stone, 2013; Hirai et al., 2018). The quasi-modes resulting from these spectra can corotate anywhere in the radial domain, making them ideal candidates for instability in both local Cartesian and global simulations.
@@
In practice, however, large-$k$ and large-$m$ modes can only be realised as long as the scales involved are above the resistive and viscous length scales. The effect of these non-ideal terms will generally be to quench the instability, particularly on smaller scales (Balbus & Hawley, 1991, 1998; Kitchatinov & Rüdiger, 2010). We plan to further investigate the effect of these diffusive terms on the MRI/SARI, especially on localised quasi-continuum SARI modes. Since the inclusion of non-ideal MHD effects breaks the up-down symmetry of MHD spectra (Goedbloed et al., 2019), it will be interesting to see how the quasi-continuum regions are modified in the structure of the spectra. We note that Legolas can analyse a number of non-ideal effects in combination, and thus can explore visco-resistive, or Hall and ambipolar effects on the ideal MHD spectrum.
@@
Magnetic field – We can conclude from Sect. 5.2 and 5.3 that a strong poloidal field inhibits large-$k$ instabilities, while a strong toroidal field inhibits large-$m>0$ instabilities: the magnetic tension produced by these small-wavelength modes would be too high (Balbus & Hawley, 1998). As a result, if the field is predominantly poloidal and strong, the MRI will be stabilised at small $k$ while negative $m$ modes can still have significant growth rates for large $k$ if $\omega_{A}$ satisfies criterion (10) (Begelman et al., 2022; Begelman & Armitage, 2023), as was also confirmed by Fig. 5. In particular, counter-rotating SARIs have one value of $m$ where the unstable range of field strengths is maximally extended, as in Fig. 7 (b). This implies that strong, poloidal fields could be unstable to extremely high-$m$ modes beyond the capabilities of our linear (and non-linear) codes: their eigenfunctions would be constrained to a tiny fraction of the domain. We also confirmed that for a purely toroidal field, a ‘complete breakdown of the most rapidly growing local modes’ (Balbus & Hawley, 1998) does indeed occur: the $k_{\text{max}}\rightarrow\infty$ for a purely toroidal field predicted by Ogilvie & Pringle (1996) quickly moves to much smaller values for even a seed poloidal field, as was noted in Fig. 6. Finally, in our exploration of the plasma-beta, we found that a dominant toroidal field has low growth rates for $\beta<1$. This was also found by Das et al. (2018) at intermediate field strengths, but they identified a new hybrid Alfvén/slow instability at stronger fields. It is unclear if a similar instability appears for our $m=10$ SARI in Case (a). According to van der Swaluw et al. (2005), these instabilities must be more MRI-like than convection-like, since the parameter range explored here corresponds to thin disks ($\sqrt{p_{1}}\leq 0.6$).
@@
Not only the fastest-growing modes are of importance for stability: the linear evolution is governed by an entire MHD spectrum of eigenmodes. We found that the outer SARI branch is usually stabilised first, implying that the quasi-continuum between the discrete SARI branches can detach from the real axis and disappear before that, or continue to exist until the inner SARIs are also stabilised. It should be noted that we did not cover SARI quasi-continua for a dominant poloidal field, but these do exist. We confirmed this for $m=1$ at weak field strength. The Appendix covers the basic theory needed for the global formulation of the eigenvalue problem, from which a local dispersion relation can be derived. We show the connection between the general dispersion relation from Blokland et al. (2005) and recent work by Das et al. (2018) and Ebrahimi & Pharr (2022). By comparing solutions to this dispersion relation with a SARI spectrum in Fig. 10, we highlighted that the local approximation is insufficient to describe outer and quasi-continuum SARI eigenmodes.
@@
N.B. is supported by an FWO Flanders fellowship titled “Magnetoseismology of accretion disks: towards the full spectrum for turbulent disks” - grant number 11J2622N. R.K. acknowledges support by the ERC Advanced Grant PROMINENT from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 833251 PROMINENT ERC-ADG 2018), the C1 project TRACESpace funded by KU Leuven, and FWO project G0B4521N.